OpenStudy (anonymous):

PLEASE HELP! Sketch the graph of the function by a) applying the Leading Coefficient Test, b) finding the zeros of the polynomial, c) plotting sufficient solution points, and d) drawing a continuous curve through the points. g(x)=x^4-4x^2

OpenStudy (jdoe0001):

a) applying the Leading Coefficient Test, g(x)= x^4-4x^2 ^ +1 b) finding the zeros of the polynomial, g(x)=x^4-4x^2 set y=0 0=x^4-4x^2 solve for "x"

OpenStudy (anonymous):

for a what is the +1 for? give me just a second to work out b!

OpenStudy (jdoe0001):

the "leading coefficient test" just means.... check the number in front of the leading term if the number is positive, the graph goes one way if the number is negative, it turns upside down

OpenStudy (anonymous):

for b i got x=0 x=2 x=-2

OpenStudy (anonymous):

is that correct?

OpenStudy (jdoe0001):

yes \(\bf y=x^4-4x^2\implies 0=x^2(x-2)(x+2)\) notice the "x squared", it means \(\bf x^2=0\implies x\cdot x=\implies x=0, \quad x=0\) is has a "multiplicity" of 2, when the root multiplicity is even, it means at that point, the graph HITS the x-axis BUT DOESN'T CROSS IT, just touches it and backs up/down

OpenStudy (jdoe0001):

now you'd just need to ... plot a couple of points like the middle x-value between -2 and 0 and 0 and 2 to see where the "hump" will be, either up or down

OpenStudy (anonymous):

how do i know which ppints to plot?

OpenStudy (jdoe0001):

well... between x-intercepts, for a quadratic... the middle point between them gives you the vertex for a quartic, like this one... it might be a bit off, but close enough so the middle or close to middle or maybe even 3, the middle value between the x-intercepts and a couple more to the side of it will do

OpenStudy (anonymous):

fwhat about (-1,-4) (-2,4)(0,0)(1,-4)(2,4)? could those be points?

OpenStudy (jdoe0001):

well... -2 is a root.... so, that'd be -2,0 more like something like .... -1.5, -1, -0.5 +1.5, +1, +0.5

OpenStudy (jdoe0001):

\(\bf y=x^4-4x^2\qquad x=\{\pm1.5, \pm1, \pm0.5\}\)

OpenStudy (anonymous):

so then those are what the points should be?? im sorry im a bit confused on how you got that

OpenStudy (jdoe0001):